Is there a formula to combine like solution concentrations to make 100%
For Instance if you have a 123% pigment solution and an 87% pigment solution, how do combine them to make a 100% solution?

Answer by stanbon(75887)

Is there a formula to combine like solution concentrations to make 100%
For Instance if you have a 123% pigment solution and an 87% pigment solution, how do combine them to make a 100% solution?
---
There is no such thing as a 123% pigment solution.
---
Let's try the math as if there was such a thing:
pigment + pigment = pigment
1.23x + 0.87y = 1.00(x+y)
Multiply thru by 100 to get:
123x + 87y = 100x+100y
23x = 23y
x = y
---
That means 1.23x + 0.87x = 1*(x+x)
2x = 2x
x = x
-----
So no matter what quantity of 123% solution you
add to 87% solution you get 100% solution.
=======================
Cheers,
Stan H.
============
Stan H.